can u help me on this problem 5(x-7)=5

5x-35=5
5x=5+35
5x=40
x=-40/5
x=-8

Found 4 solutions by longjonsilver, Dali, zynan88, Alan3354:
Answer by longjonsilver(2297)

5(x-7)=5 -- divide both sides by 5
5(x-7)/5 = 5/5

x-7 = 1 -- add 7 to both sides
x-7+7 = 1+7

--> x = 8

jon

Answer by Dali(1)

5x-35=5
+35 +35
5x=40
x=8

Answer by zynan88(4)

5x-35=5
5x=5+35
5x=40
x=-40/5
x=-8

Answer by Alan3354(69443)

5x-35=5
Divide by 5
x-7 = 1
Add 7
x = 8